OpenStudy (anonymous):

The table below shows two equations: Equation 1 |2x - 3| + 5 = 4 Equation 2 |5x + 3| - 10 = 3

OpenStudy (anonymous):

Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution, and equation 2 has solutions x = 2, -3.2. The solutions to equation 1 are x = 1, 2, and equation 2 has no solution. The solutions to equation 1 are x = 1, 2, and equation 2 has solutions x = 2, -3.2.

OpenStudy (anonymous):

I think the answer to this one is the second option.

OpenStudy (anonymous):

cw gots you

OpenStudy (cwrw238):

equation 1 has no solutions because |2x-3| = -1 and the absolute value must be positive

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

second option